Hayhayz:

The functions q and r are defined as follows. q(x)=2x-1 r(x)=-2x^2+2 Find the value of q(r(−3))=

Hayhayz:

@HuskyNation @JustSaiyan @Justin234 @Elsa213 can anyone halp? lol

Elsa213:

Hayzz I em heree Pewpewpew lemme see e.e

Hayhayz:

okeyyy com to da rescue for meh pls <3

Elsa213:

are der answer choices? e.e

Hayhayz:

nupe

Elsa213:

Err mehh gawdd das a tough one dennn

Hayhayz:

aahhh

Elsa213:

Wat answer deed yew get? e.e

Hayhayz:

39 somehow lol bu its wrong

Elsa213:

Hrmmm I solved et and got du same answer o:

Elsa213:

@HuskyNation

babegurl2334:

wait why does one punch man still go to school?

Hayhayz:

shhhh

babegurl2334:

lol one punch didnt pass the first grade WEXD

Hayhayz:

stop making my self esteem vanish gurl

babegurl2334:

most of the times people sound like pedos and call me bae but i am proud of you

random231:

first put -3 in place of x in function r(x)

random231:



random231:

the value you get from this, you put in place of x in q(x)

MARC:

working : \(=q[(r(-3)]\) \(=q[-2(\color{red}{-3})^2+2)]\) \(=q[-2(9)+2]\) \(=q(-18+2)\) \(=q(-16)\) \(=2(\color{blue}{-16})-1\) \(=-32-1\) \(=?\) solve it! @Hayhayz

mathmale:

Hayhayz: Has this helped you? If you're done, please close this post. Otherwise comment or ask questions that might take you to completion.

Marziman:

Hayzio just for the future, when you have a question like this where it combines q(x) and r(x), you just combine the functions. The X was replaced by -3 in r(x) and q(x) so in the functions you simply replace those X variables with -3. Ez Pz